# web_appendix

### v4

@@ 7 Application
We use the yearly 1966-to-2008 panel dataset of Li et al. (2014), covering 48 US states (Alaska and Hawaii are excluded). In view of the long duration of this panel, it is important to keep in mind that our estimators only assume parallel trends across pairs of consecutive years, not over the panel’s entire duration. For each state×$\times$×year cell $(i,t)$, the data contains $Z_{i,t}$, the total (state plus federal) gasoline tax in cents per gallon, $D_{i,t}$, the log tax-inclusive price of gasoline, and $Y_{i,t}$, the log gasoline consumption per adult. Our goal is to estimate the effect of gasoline taxes on gasoline consumption and prices, and to estimate the price-elasticity of gasoline consumption, using taxes as an instrument. Instead, Li et al. (2014) jointly estimate the effect of gasoline taxes and tax-exclusive prices on consumption, using a TWFE regression with two treatments. Between each pair of consecutive periods, the tax-exclusive price changes in all states, so this treatment does not have stayers and its effect cannot be estimated using the estimators proposed in this paper. Thus, our estimates cannot be compared to those of Li et al. (2014).
Switching cells, and how they compare to the entire sample.
@@
Let $\mathcal{S}$ be the set of switching $(i,t)$ cells such that $Z_{i,t}\neq Z_{i,t-1}$ but $Z_{i^{\prime},t}=Z_{i^{\prime},t-1}$ for some $i^{\prime}$. The second condition drops from the estimation seven pairs of consecutive time periods between which the federal gasoline tax changed, thus implying that all states experienced a change of their tax. $\mathcal{S}$ includes 384 cells, so effects of taxes on gasoline prices and consumptions can be estimated for 19% of the 2,016 state×$\times$×year cells for which $Z_{i,t}-Z_{i,t-1}$ can be computed. Table 1 below compares some observable characteristics of switchers and stayers. Switchers seem slightly over-represented in the later years of the panel: $t$ is on average 2.5 years larger for switchers than for stayers, and the difference is significant. On the other hand, switchers are not more populated than stayers, and their gasoline consumption and gasoline price in 1966 are not significantly different from that of stayers. Thus, there is no strong indication that the cells in $\mathcal{S}$ are a very selected subgroup.
Table 1: Comparing switchers and stayers
Dependent Variables: $t$ Adult Population log(quantity)1966 log(price)1966
Constant 1,986.7 3,691,608.0 -0.5161 3.471
(0.2739) (577,164.0) (0.0210) (0.0054)
$\mathbf{1}\{Z_{i,t}\neq Z_{i,t-1}\}$ 2.481 39,588.0 -0.0099 0.0014
(0.7519) (320,342.1) (0.0096) (0.0029)
N 2,016 2,016 2,016 2,016
Notes: The table show the results of regressions of some dependent variables on a constant and an indicator for switching cells. The standard errors shown in parentheses are clustered at the state level.
Distribution of taxes.
@@
As an example, the top panel of Figure 1 below shows the distribution of $Z_{g,1987}$ for 1987-to-1988 stayers, while the bottom panel shows the distribution for 1987-to-1988 switchers. The figure shows that there are many values of $Z_{g,1987}$ such that only one or two states have that value, so $Z_{g,1987}$ is close to being continuously distributed. Moreover, all switchers $g$ are such that
$\min_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1987}}Z_{g^{\prime},1987}\leq Z_{g,1987}\leq\max_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1988}}Z_{g^{\prime},1987}.$
@@
Thus, Assumption 4 seems to hold for this pair of years. $(1987,1988)$ is not atypical. While $Z_{i,t}$ varies less across states in the first years of the panel, there are many other years where $Z_{i,t}$ is close to being continuously distributed. Similarly, almost 95% of cells in $\mathcal{S}$ are such that $\min_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}\leq Z_{i,t-1}\leq\max_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}$. Dropping the few cells that do not satisfy this condition barely changes the results presented below.
@@
Figure 2 below shows the distribution of $Z_{i,t}-Z_{i,t-1}$ for the 384 cells in $\mathcal{S}$. The majority experience an increase in their taxes, but 38 cells experience a decrease. The average value of $|Z_{i,t}-Z_{i,t-1}|$ is equal to 1.61 cents, while prior to the tax change, switchers’ average gasoline price is equal to 112 cents: our estimators leverage small changes in taxes relative to gasoline prices. Finally, $\min_{(i,t)\in\mathcal{S}}|Z_{i,t}-Z_{i,t-1}|=0.05$ : some switchers experience a very small change in their taxes.
Figure 2: Distribution of tax changes between consecutive periods
Reduced-form and first-stage AS and WAS estimates.
@@
Table 2 below shows the AS and doubly-robust WAS estimates of the reduced-form (Panel A) and first-stage (Panel B) effects of taxes on quantities and prices. We follow results from Section 5, and estimate the reduced-form and the first-stage controlling for lagged prices $D_{t-1}$, to ensure that the resulting IV estimator is robust to heterogeneous effects over time. Reduced-form and first-stage estimators where $D_{t-1}$ is not controlled for are not very different, but controlling for $D_{t-1}$ reduces the standard error of the first-stage estimator. In Column (1), the estimators are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$. In Column (2), a polynomial of order 2 is used in those estimations. 10-folds cross-validation selects a polynomial of order two for $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and a polynomial of order one for all the other conditional expectations. Thus, polynomials of order 1 and 2 are in line with those selected by cross validation. Standard errors clustered at the state level, computed following (18) and (19), are shown below the estimates, between parentheses. All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers. Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal. In Panel A Column (1), the AS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.55 percent on average for the switchers. That effect is significant at the 5% level, but it becomes smaller and insignificant when one uses a quadratic model to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$. The WAS estimates are slightly lower than, but close to, the AS estimates, and they are significant irrespective of the polynomial order used in the estimation. As predicted by Proposition 1, the standard errors of the WAS estimators are almost 3 times smaller than that of the AS estimators. Equality tests that the AS and WAS effects are equal are not rejected. In Panel B, the AS estimates of the first-stage effect are insignificant. Importantly, this implies that an IV-AS estimator of the price elasticity of gasoline consumption cannot be used: this estimator does not have a significant first stage. The WAS estimates are significant, and they indicate that if gasoline tax increases by 1 cent on average, prices increase by around 0.5 percent on average for the switchers. Again, the differences between the AS and WAS effects of taxes on prices are insignificant.
Table 2: Effects of gasoline tax on quantities consumed and prices
(1) Linear model (2) Quadratic model
Panel A: Reduced-form effect of taxes on quantities consumed
AS -0.0055 -0.0034
(0.0027) (0.0032)
WAS -0.0038 -0.0034
(0.0010) (0.0011)
Observations 1,632 1,632
P-value 0.4482 0.9974
Panel B: First-stage effect of taxes on prices
AS 0.0042 0.0047
(0.0024) (0.0025)
WAS 0.0056 0.0056
(0.0009) (0.0008)
Observations 1,632 1,632
P-value 0.4729 0.6798
@@
Notes: All estimators in the table are computed using the data of Li et al. (2014). Panel A (resp. B) shows the AS and doubly-robust WAS estimates of the reduced-form (resp. first-stage) effect of taxes on quantities (resp. prices). All estimates control for the lag of prices. In Column (1), estimates are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$ and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$. In Column (2), estimates are computed using a polynomial of order 2 in those estimations. Standard errors clustered at the state level, computed following (18) and (19) are shown below the estimates, between parentheses. All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers. Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
Placebo analysis.
@@
Table 3 below shows placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects. The placebo estimators are analogous to the actual estimators, but they replace $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and they restrict the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$. The placebo WAS estimates are small and insignificant, both for quantities and prices. The placebo AS estimates are larger for quantities, but they are insignificant, and less precisely estimated. This placebo analysis shows that before switchers change their gasoline taxes, switchers’ and stayers’ consumption of gasoline and gasoline prices do not follow detectably different evolutions. As a robustness check, we reestimate the AS and WAS in the placebo subsample, to ensure that effects are estimated in a subsample for which the identifying assumption can be placebo tested, and also because in that subsample estimators remain valid if the first lag of taxes affect current gasoline prices and quantities. WAS reduced-form effects are very close to those in Table 2. WAS first-stage effects are 25 to 35% smaller, though they are still positive and highly significant.
Table 3: Placebo effects of gasoline tax on quantities consumed and prices
(1) Linear model (2) Quadratic model
Panel A: Reduced-form placebo effect of taxes on quantities consumed
AS 0.0039 0.0055
(0.0035) (0.0036)
WAS 0.0001 0.0012
(0.0017) (0.0017)
Observations 1,059 1,059
Panel B: First-stage placebo effect of taxes on prices
AS 0.0006 0.0009
(0.0056) (0.0053)
WAS 0.0014 0.0013
(0.0017) (0.0015)
Observations 1,059 1,059
@@
Notes: The table shows the placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects of taxes on quantities and prices. The estimators and their standard errors are computed as the actual estimators, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$.
IV-WAS estimate of the price-elasticity of gasoline consumption.
@@
Table 4 shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption. As the instrument’s first stage is not very strong and the sample effectively only has 48 observations, asymptotic approximations may not be reliable for inference. In line with that conjecture, we find that the bootstrap distributions of the three estimators in Table 4 are non-normal, with some outliers. Therefore, we use the percentile bootstrap for inference, clustering the bootstrap at the state level. Reassuringly, these confidence intervals have nominal coverage in simulations tailored to our application. [^7] Here is the DGP used in our simulations. We estimate TWFE regressions of $Y_{i,t}$ on state and year fixed effects and $Z_{i,t}$, and of $D_{i,t}$ on state and year fixed effects and $Z_{i,t}$. We let $\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y}_{i,t}$ and $\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D}_{i,t}$ denote the resulting regression decompositions. In each simulation, the simulated instrument is just the actual instrument, while the simulated outcomes and treatments are respectively equal to $Y^{s}_{i,t}=\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y,s}_{i,t},$ and $D^{s}_{i,t}=\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D,s}_{i,t},$ where the vector of simulated residuals $(\epsilon^{Y,s}_{g,1},...,\epsilon^{Y,s}_{g,T},\epsilon^{D,s}_{g,1},...,\epsilon^{D,s}_{g,T})$ is drawn at random and with replacement from the estimated vectors of residuals $((\epsilon^{Y}_{g^{\prime},1},...,\epsilon^{Y}_{g^{\prime},T},\epsilon^{D}_{g^{\prime},1},...,\epsilon^{D}_{g^{\prime},T}))_{g^{\prime}\in\{1,...,G\}}$. Thus, the first-stage and reduced-form effects, the correlation between the reduced-form and first-stage residuals, and the residuals’ serial correlation are the same as in the sample. The IV-WAS estimates are negative, significant, and larger than -1, though their confidence intervals contain -1. We compare those estimates to a 2SLS-TWFE estimator, computed via a 2SLS regression of $Y_{i,t}$ on $D_{i,t}$ and state and year fixed effects, using $Z_{i,t}$ as the instrument. The 2SLS-TWFE coefficient is equal to -1.0836 (bootstrap confidence interval=$[-2.1207,-0.4405]$), which is 60% larger in absolute value than the IV-WAS estimate in Column (1), and almost 80% larger than that in Column (2), though the 2SLS-TWFE coefficient does not significantly differ from the two IV-WAS estimates (P-value = 0.320 and 0.232 , respectively). Interestingly, the confidence interval of the 2SLS-TWFE coefficient is almost 80% wider than that of the IV-WAS coefficient in Column (1) and 27% wider than that in Column (2), thus showing that using a more robust estimator does not always come with a substantive precision cost.
Table 4: IV estimators of the price-elasticity of gasoline consumption
(1) Linear model (2) Quadratic model
IV-WAS -0.6773 -0.6130
[-1.2101,-0.2622] [-1.3183,-0.0004]
Observations 1,632 1,632
@@
Notes: The table shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption, computed using the data of Li et al. (2014). Bootstrap confidence intervals are shown below the estimates. They are computed with 500 bootstrap replications, clustered at the state level.

### v5

@@ 6.1 Data and design
We use the yearly 1966-to-2008 panel dataset of Li et al. (2014), covering 48 US states (Alaska and Hawaii are excluded). This is a long panel, but recall that our estimators assume parallel trends only across pairs of consecutive years. For each state×$\times$×year cell $(i,t)$, the data contains $Z_{i,t}$, the total (state plus federal) gasoline tax in cents per gallon, $D_{i,t}$, the log tax-inclusive price of gasoline, and $Y_{i,t}$, the log gasoline consumption per adult. We seek to estimate the effect of taxes on gasoline consumption and prices, and the price-elasticity of gasoline consumption using taxes as an instrument. [^8] Instead, Li et al. (2014) jointly estimate the effect of gasoline taxes and tax-exclusive prices on consumption, using a TWFE regression with two treatments. Between each pair of consecutive periods, the tax-exclusive price changes in all states, so this treatment does not have stayers and its effect cannot be estimated using the estimators proposed in this paper. Thus, our estimates cannot be compared to those of Li et al. (2014).
@@
Let $\mathcal{S}$ be the set of switching $(i,t)$ cells such that $Z_{i,t}\neq Z_{i,t-1}$ but $Z_{i^{\prime},t}=Z_{i^{\prime},t-1}$ for some $i^{\prime}$. The second condition drops from the estimation seven pairs of consecutive time periods between which the federal gasoline tax changed, thus implying that all states experienced a change of their tax. $\mathcal{S}$ includes 384 cells, namely 19% of the 2,016 state×$\times$×year cells for which $Z_{i,t}-Z_{i,t-1}$ can be computed. Table LABEL:table:balancing in the web appendix shows that cells in $\mathcal{S}$ are fairly similar to other cells on a number of observable characteristics. Thus, there is no strong indication that the effects we estimate below apply to a very selected subgroup.
@@
As an example, the top panel of Figure 1 below shows the distribution of $Z_{i,1987}$ for 1987-to-1988 stayers, while the bottom panel shows the distribution for 1987-to-1988 switchers. The figure shows that there are many values of $Z_{i,1987}$ such that only one or two states have that value, so $Z_{i,1987}$ is close to being continuously distributed. Moreover, all switchers $i$ are such that
$\min_{i^{\prime}:Z_{i^{\prime},1988}=Z_{i^{\prime},1987}}Z_{i^{\prime},1987}\leq Z_{i,1987}\leq\max_{i^{\prime}:Z_{i^{\prime},1988}=Z_{i^{\prime},1988}}Z_{i^{\prime},1987}.$
@@
Thus, Assumption 4 seems to hold for this pair of years. $(1987,1988)$ is not atypical. While $Z_{i,t}$ varies less across states in the first years of the panel, there are many other years where $Z_{i,t}$ is close to being continuously distributed. Similarly, almost 95% of cells in $\mathcal{S}$ are such that $\min_{i^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}\leq Z_{i,t-1}\leq\max_{i^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}$. Dropping the few cells that do not satisfy this condition barely changes the results presented below.
@@
Turning to the distribution of the tax changes $Z_{i,t}-Z_{i,t-1}$, 90% of the cells in $\mathcal{S}$ experience an increase in their taxes, and 10% experience a decrease. The average value of $|Z_{i,t}-Z_{i,t-1}|$ is equal to 1.61 cents, while prior to the tax change, switchers’ average gasoline price is equal to 112 cents: our estimators leverage small changes in taxes relative to gasoline prices. Finally, $\min_{(i,t)\in\mathcal{S}}|Z_{i,t}-Z_{i,t-1}|=0.05$ : some switchers experience a very small change in their taxes.
6.2 Reduced-form and first-stage estimates
@@
Table 6.2 below shows the AS and WAS estimates of the reduced-form (Panel A) and first-stage (Panel B) effects of taxes on gasoline quantities and prices. We control for lagged prices $D_{t-1}$, to ensure that the resulting IV estimate is robust to heterogeneous effects over time, as explained in Section 5.1. Estimates are computed with cross-fitting, using 10 splits. We use a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E[S/\Delta D|Z_{t-1},D_{t-1}]$, $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$. 10-folds cross-validation selects a polynomial of order one for all those conditional expectations, except for $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$ where it selects a polynomial of order two. Thus, polynomials of order one are in line with those selected by cross validation. [^9] In a previous version of this paper (de Chaisemartin et al., 2022), we had shown results very similar to those below with polynomials of order two and without cross fitting. With polynomials of order two and cross fitting, estimates become noisy: due to the small sample size, $\widehat{P}(S_{t}=0|Z_{t-1},D_{t-1})$ is close to zero for some $(i,t)$. Standard errors clustered at the state level, computed following (LABEL:eq:ASIFT>=3) and (LABEL:eq:WASIFT>=3) in the web appendix, are shown next to the estimates, between parentheses. Estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods are excluded as they do not have stayers. The last line of each panel shows the p-value of a test that AS=WAS.
In Panel A, the AS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.43 percent on average for the switchers. That effect is insignificant. The WAS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.36 percent, an effect that is significant. In Panel B, the AS estimate of the first-stage effect is positive and insignificant. The WAS estimate is significant, and indicates that if gasoline tax increases by 1 cent on average, prices increase by 0.58 percent. In both panels, the AS estimates are close to, and not significantly different from, the WAS estimates, but they are insignificant because they are markedly less precise: their standard errors are almost three times larger, as predicted by Proposition 1. Therefore, in what follows we focus on the WAS estimates.
Table 1: Effects of gasoline tax on quantities and prices
@@
Notes: Panel A (resp. B) shows the AS and WAS estimates of the reduced-form (resp. first-stage) effect of taxes on gasoline quantities (resp. prices). All estimates control for the lag of prices. All conditional expectations are estimated using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$, with cross-fitting, using ten splits. Standard errors clustered at the state level are shown next to the estimates, between parentheses. The last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
6.3 Placebo analysis
@@
Table LABEL:table:placeboRFandFS below shows “pre-trends” or “placebo” WAS estimates of the reduced-form and first-stage effects. The placebos are analogous to the actual estimates, but they replace $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and they restrict the sample to states whose taxes did not change between $t-2$ and $t-1$. This reduces the sample size by around a third. The placebos are small and insignificant, both for quantities and prices. This shows that before switchers change their gasoline taxes, switchers’ and stayers’ consumption and prices do not follow detectably different evolutions. In Table LABEL:table:RFandFS_placebosample of the web appendix we reestimate the WAS in the placebo subsample. In that subsample, the WAS remains valid if the first lag of taxes affect current gasoline prices and quantities, as discussed in Section 5.2. The first-stage and reduced-form estimates therein are close to those in Table 6.2.
@@
Notes: The table shows placebo WAS estimates of the reduced-form and first-stage effects of taxes on quantities and prices. The estimates and their standard errors are computed as the actual estimates, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample to states whose taxes did not change between $t-2$ and $t-1$.
6.4 IV estimate
@@
Table LABEL:table:IV shows an IV-WAS estimate of the price-elasticity of gasoline consumption. As the instrument’s first stage is not very strong and the sample only has 48 observations, asymptotic approximations may not be reliable for inference. In line with that conjecture, we find that the bootstrap distribution of the estimator in Table LABEL:table:IV is non-normal, with some outliers. Therefore, for inference we use a percentile bootstrap confidence interval, clustered at the state level. Reassuringly, that confidence interval has close to nominal coverage in simulations tailored to our application. [^10] Here is the DGP used in our simulations. We estimate TWFE regressions of $Y_{i,t}$ on state and year fixed effects and $Z_{i,t}$, and of $D_{i,t}$ on state and year fixed effects and $Z_{i,t}$. We let $\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y}_{i,t}$ and $\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D}_{i,t}$ denote the resulting regression decompositions. In each simulation, the simulated instrument is just the actual instrument, while the simulated outcomes and treatments are respectively equal to $Y^{s}_{i,t}=\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y,s}_{i,t},$ and $D^{s}_{i,t}=\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D,s}_{i,t},$ where the vector of simulated residuals $(\epsilon^{Y,s}_{i,1},...,\epsilon^{Y,s}_{i,T},\epsilon^{D,s}_{i,1},...,\epsilon^{D,s}_{i,T})$ is drawn at random and with replacement from the estimated vectors of residuals $((\epsilon^{Y}_{i^{\prime},1},...,\epsilon^{Y}_{i^{\prime},T},\epsilon^{D}_{i^{\prime},1},...,\epsilon^{D}_{i^{\prime},T}))_{i^{\prime}\in\{1,...,n\}}$. Thus, the first-stage and reduced-form effects, the correlation between the reduced-form and first-stage residuals, and the residuals’ serial correlation are the same as in the sample. The IV-WAS estimate is negative and significant at any conventional level. It is more negative than, but not significantly different from, estimates reported in recent literature (Coglianese et al., 2017).
@@
Notes: The table shows an IV-WAS estimate of the price-elasticity of gasoline consumption, controlling for the lag of prices. All conditional expectations are estimated using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$, with cross-fitting, using ten splits. A percentile-bootstrap confidence interval, computed with 500 bootstrap replications clustered at the state level, is shown below the estimate.